OpenStudy (jackfrost33):

how do you express 2a^-4/ 5b^-2 with positive exponents?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (jackfrost33):

nope

OpenStudy (skullpatrol):

$$\Huge \dfrac{2a^{-4}}{5b^{-2}}$$

OpenStudy (jackfrost33):

this is a new concept and i dont get it at all my lesson does not really teach you how to do this

OpenStudy (skullpatrol):

Use the definition of a negative exponent.

OpenStudy (jackfrost33):

okay

OpenStudy (skullpatrol):

$$ \Huge a^{-n} = \dfrac{1}{a^n}$$

OpenStudy (skullpatrol):

$$\Huge \dfrac{2a^{-4}}{5b^{-2}}$$ $$\Huge \dfrac{2/a^{4}}{5/b^{2}}$$

satellite73 (satellite73):

you got this?

satellite73 (satellite73):

no one wants to see a compound fraction if the exponent is negative with a term in the numerator (top) bring it down if the exponent is negative with a term in the deminator, bring it up

OpenStudy (skullpatrol):

$$\Huge \dfrac{\dfrac{2}{a^4}}{\dfrac{5}{b^2}}$$

OpenStudy (skullpatrol):

$$\Huge \dfrac{\dfrac{2}{a^4} • a^4 b^2}{\dfrac{5}{b^2} • a^4 b^2}$$